An airliner weighing 12,905 kg needs to achieve a speed of 84.6 m/s to take off. If the length of the runway is 1,071 m, and the engines develop a constant force, what is the force necessary to achieve the necessary speed to take off? Enter your answer as the number of kiloNewtons (kN) required.
1
Expert's answer
2020-11-16T07:47:48-0500

"s=v^2/(2a)\to a=v^2/(2s)=84.6^2/(2\cdot1071)=3.34(m/s^2)"


"F=ma=12905\cdot3.34=43120(N)=43.12(kN)". Answer